# LDP for invariant measures of the Allen-Cahn equation with inhomogeneous boundary conditions and multiplicative noise

### v1

@@ 4.1 Exponential estimate
$\mathbb{P}\left(\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}\geq\bar{R}_{r}\right).$
If we show that there exist $C_{1},C_{2}>0$ such that
$\mathbb{E}\exp\left(\frac{(\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}-C_{1})^{2}}{\epsilon}\right)\leq C_{2},$
then, thanks to Chebyshev inequality, we immediately obtain that there exists $\bar{R}_{r}>C_{1}$ and $\epsilon_{r}$ depends on $r$ such that
$\mathbb{P}\left(\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}\geq\bar{R}_{r}\right)\leq\exp(-\frac{r}{\epsilon}),\epsilon<\epsilon_{r}.$ (4.14)
According to standard estimates for equation (1.3), as in the proof of Proposition 6.1 of [cerrai2003stochastic],
$\begin{split}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}&\leq C\left(\epsilon^{\frac{1}{2}}\sup_{s\leq t}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{E}+\lambda^{\frac{1}{3}}\sup_{s\leq t}\|\bar{u}_{\epsilon}^{0}(s)\|_{E}^{\frac{1}{3}}+1\right)\\ &\leq C\left(\epsilon^{\frac{1}{2}}\sup_{s\leq t}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{E}+\lambda^{\frac{1}{2}}+1\right)+\frac{1}{2}\sup_{s\leq t}\|\bar{u}_{\epsilon}^{0}(s)\|_{E},\end{split}$ (4.15)
where the last inequality is due to Young’s inequality. Then,
$\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}\leq C\left(\epsilon^{\frac{1}{2}}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{E}+\lambda^{\frac{1}{2}}+1\right).$
Thus,
$\frac{(\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}-C(1+\lambda^{\frac{1}{2}}))^{2}}{\epsilon}\leq C^{2}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{E}^{2}$
and
$\ \ \ \ \mathbb{E}\exp{\left(\frac{(\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}-C(1+\lambda^{\frac{1}{2}}))^{2}}{\epsilon}\right)}$
$=\sum_{i\in\mathbb{N}^{0}}\frac{1}{i!}\mathbb{E}\left(\frac{(\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}-C(1+\lambda^{\frac{1}{2}}))^{2}}{\epsilon}\right)^{i}$
$\leq\sum_{i\in\mathbb{N}^{0}}\frac{C^{2i}}{i!}\mathbb{E}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{E}^{2i}.$
Again, for $p\geq p^{\star}$ while $k^{\star},\alpha$ remain the same, we apply the same computation as (4.9). Then,
$\mathbb{E}\left[\sup_{s\leq t}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{W^{k^{\star},p}}^{p}\right]\leq C_{\lambda,p}^{p}\mathbb{E}\left[\sup_{s\leq t}\|\Gamma^{\alpha}(\bar{u}_{\epsilon}^{0})(s)\|_{L^{p}}^{p}\right].$ (4.16)
Here, by Proposition 4.5 of [cerrai2003stochastic]
$C_{\lambda,p}\leq(\frac{2}{\lambda p})^{\frac{1}{p}}\cdot(\frac{1}{1-(2\alpha+\frac{1}{2})}+\frac{2}{\lambda})^{\frac{1}{2}}\cdot\left(\frac{1}{1-\frac{p(1+k^{\star}/2-\alpha)}{p-1}}+\frac{2(p-1)}{p\lambda}\right)^{\frac{p-1}{p}},$
and $\lim_{\lambda\rightarrow\infty}C_{\lambda,p}=0$. Note, for any $t\geq 0$, $\xi\in[-L,L]$
$\mathbb{E}|\Gamma^{\alpha}(\bar{u}_{\epsilon}^{0})(t,\xi)|^{2}$ $\leq\sum_{k\in\mathbb{N}}\int_{0}^{t}s^{-2\alpha}e^{-2(\frac{L^{-1}k}{2})^{2}s}\mathbb{E}(1+\|\bar{u}_{\epsilon}^{0}(s)\|_{E})^{2}\;ds\cdot e_{k}^{2}(\xi)$
$\leq C\sum_{k\in\mathbb{N}}\frac{1}{(L^{-1}k\pi)^{2(1-2\alpha)}}\mathbb{E}\sup_{s\leq t}(1+\|\bar{u}_{\epsilon}^{0}(s)\|_{E})^{2}$
$\leq C_{\alpha}\mathbb{E}\sup_{s\leq t}(1+\|\bar{u}_{\epsilon}^{0}(s)\|_{E})^{2},$
where $C_{\alpha}>0$ depends on $\alpha$ and does not depend on $t$. Therefore, by the Burkholder–Davis–Gundy inequality there exists $C>0$, such that
$\mathbb{E}\left[\sup_{s\leq t}\|\Gamma^{\alpha}(\bar{u}_{\epsilon}^{0})(s)\|_{L^{p}}^{p}\right]$ $\leq\mathbb{E}\int_{-L}^{L}\sup_{s\leq t}|\Gamma^{\alpha}(\bar{u}_{\epsilon}^{0})(s,\xi)|^{p}\;d\xi$
$\leq\int_{-L}^{L}(\mathbb{E}|\Gamma^{\alpha}(\bar{u}_{\epsilon}^{0})(t,\xi)|^{2})^{\frac{p}{2}}\;d\xi$
$\leq C^{p}C_{\alpha}^{\frac{p}{2}}(1+\mathbb{E}\sup_{s\leq t}\|\bar{u}_{\epsilon}^{0}(s)\|_{E}^{p}).$
Since $C_{\lambda,p},C_{\alpha}$ do not depend on $t$,
$\mathbb{E}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{E}^{p}\leq\mathbb{E}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p}}^{p}\leq C^{p}C_{\alpha}^{\frac{p}{2}}C_{\lambda,p}^{p}(1+\mathbb{E}\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}^{p}).$ (4.17)
We choose $\lambda$ sufficiently large such that $C\cdot C_{\alpha}^{\frac{1}{2}}C_{\lambda,p}\leq\frac{1}{2}$ and substitute (4.17) into (4.15). Then,
$\mathbb{E}\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}^{p}\leq\frac{1}{4}\mathbb{E}\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}^{p}+\epsilon^{\frac{p}{2}}C^{p}C_{\alpha}^{\frac{p}{2}}C_{\lambda,p}^{p}(1+\mathbb{E}\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}^{p})+(C+\lambda^{\frac{1}{2}})^{p}.$
This implies that for $p\geq p^{\star}$, there exists $C>0$ independent of $p$ such that
$\mathbb{E}\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}^{p}\leq 2(C+\lambda^{\frac{1}{2}})^{p}.$ (4.18)
Moreover, by Jensen’s inequality, for any $0\leq p\leq p^{\star}$
$\mathbb{E}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{E}^{p}$ $\leq\mathbb{E}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(t)\|_{W^{k^{\star},p^{\star}}}^{p}$
$\leq\left(2^{p^{\star}}C_{\alpha}^{\frac{p^{\star}}{2}}(1+\mathbb{E}\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}^{p^{\star}})\right)^{\frac{p}{p^{\star}}}$
$\leq C,$
where the last inequality is due to (4.18). Finally, for $\lambda$ sufficiently large, we have
$\ \ \ \ \mathbb{E}\exp{\left(\frac{(\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}-C(1+\lambda^{\frac{1}{2}}))^{2}}{\epsilon}\right)}$
$\leq\sum_{i\in\mathbb{N}^{0}}\frac{C^{2i}}{i!}\mathbb{E}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{E}^{2i}$
$=\sum_{i=0}^{[p^{\star}/2]}\frac{C^{2i}}{i!}\mathbb{E}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{E}^{2i}+\sum^{\infty}_{i=[p^{\star}/2]+1}\frac{C^{2i}}{i!}\mathbb{E}\sup_{t\geq 0}\|\gamma_{\lambda}(\bar{u}_{\epsilon}^{0})(s)\|_{E}^{2i}$
$\leq Cp^{\star}+\sum^{\infty}_{i=[p^{\star}/2]+1}\frac{C^{4i}}{i!}(1+\mathbb{E}\sup_{t\geq 0}\|\bar{u}_{\epsilon}^{0}(t)\|_{E}^{2i})$
$\leq Cp^{\star}+\sum^{\infty}_{i=[p^{\star}/2]+1}\frac{C^{4i}}{i!}(C+\lambda^{\frac{1}{2}})^{2i}$
$\leq Cp^{\star}+e^{C^{4}(C+\lambda^{\frac{1}{2}})^{2}}\leq C_{\lambda,k^{\star},p^{\star},\alpha},$
where $C_{\lambda,k^{\star},p^{\star},\alpha}>0$ is some constant that depends on $k^{\star}$, $p^{\star}$, $\alpha$ and $\lambda$. This yields (4.14). Combined with (4.13), one derives (4.8) and completes the proof. ∎
4.2 Lower bound
@@
Fix $\zeta\in E$, $\delta>0$, $\gamma>0$ and $T>0$. We assume that $U(\zeta)<\infty$ otherwise the result would be trivial. We set $\left\{z^{x}\right\}_{x\in E}\subset C([0,T];E)$ to be a family of paths satisfying
@@
Moreover, acccording to Section 2, $\left\{\mathcal{L}(\bar{u}_{\epsilon}^{x})\right\}_{\epsilon>0}$ satisfies the Freidlin-Wentzell uniform large deviations principle in $C([0,T];E)$. So, for every $r_{0},\gamma>0$ there exists $\epsilon_{0}>0$ such that $\gamma/2+\epsilon_{0}\ln\frac{1}{2}>0$ and for any $z^{x}\in C([0,T];E)$ with $I_{T}^{x}(z^{x})\leq r_{0}$,
$\inf_{x\in B_{E}(0,R)}\mathbb{P}\left(\left\|\bar{u}_{\epsilon}^{x}-z^{x}\right\|_{C([0,T];E)}<\delta/2\right)\geq\inf_{x\in B_{E}(0,R)}\exp\left(-\frac{1}{\epsilon}\left[I_{T}^{x}(z^{x})+\gamma/2+\epsilon\ln\frac{1}{2}\right]\right)$ (4.21)
for every $\epsilon\leq\epsilon_{0}$. Since the family $\left\{\bar{\mu}_{\epsilon}\right\}_{\epsilon>0}$ is the invariant measures of equation (1.3), from (4.19), (4.20) and (4.21), we have
@@
For this, denote the solution of skeleton function equation (2.3) as $z^{x}_{f}$. By the definition of $U(\zeta)$, there exist $T_{2}>0$, $z_{\bar{f}}^{m_{L}-\psi}\in C([0,T];E)$ and $\bar{f}\in L^{2}([0,T];H)$ such that $z^{m_{L}-\psi}_{\bar{f}}(0)=m_{L}-\psi$, $z^{m_{L}-\psi}_{\bar{f}}(T_{2})=\zeta$ and
@@
Thus, by the compactness of $\overline{B_{W^{k^{\star},p^{\star}}}(0,R)}$ in $E$ and (2.8), for any $\lambda>0$ we can take $T_{1}=T_{1}(\lambda)$ large enough such that
$\sup_{x\in\overline{B_{W^{k^{\star},p^{\star}}}(0,R)}}\left\|z^{x}_{0}(T_{1})-(m_{L}-\psi)\right\|_{E}\leq\lambda.$
@@
$\sup_{x\in\overline{B_{W^{k^{\star},p^{\star}}}(0,R)}}\left\|z^{x}_{f}(T)-\zeta\right\|_{E}$ $\leq\sup_{y\in B_{E}(0,\lambda)}\left\|z^{y}_{\bar{f}}(T_{2})-z^{0}_{\bar{f}}(T_{2})\right\|_{E}$
$\leq\sup_{y\in B_{E}(0,\lambda)}c(T,\|y\|_{E},U(\zeta))\cdot\|y\|_{E}.$
@@
In conclusion, the paths $z^{x}(t)=z^{x}_{f}(t)$ satisfies the two desired conditions. The proof is complete.
4.3 Upper bound

### v2

@@ 4.1 Exponential estimates
$\lim_{T\rightarrow\infty}\frac{1}{T}\int_{0}^{T}\mathbb{P}\left(\sup_{s\in[0,t]}\|\bar{u}_{\epsilon}^{0}(s)\|_{E}\geq\bar{R}_{r}\right)\;dt.$
Notice that by (3.12), the $C([0,t];E)$ norm of $z_{f}^{0}$ is bounded by $\|f\|_{L^{2}(0,t;H)}$ for all $t\geq 0$. Then, for any $r>0$, the level set $K^{0}_{t}$ is bounded in $C([0,t];E)$ uniformly with respect to all $t\geq 0$. Therefore, suppose we can prove that for any $r>0$ and all $t\geq 0$, there exists $\epsilon_{r}>0$ that does not depend on $t$, such that for any $\epsilon<\epsilon_{r}$
$\mathbb{P}\left(\text{dist}_{C([0,t];E)}(\bar{u}_{\epsilon}^{0},K^{0}_{t}(2r))\geq M\right)\leq\exp({-\frac{r}{\epsilon}})$
holds, then the proof is complete. Here, $M>0$ depends on $r$ and will be chosen later. To show that, we define functions from $C([0,t];E)$ to $\mathbb{R}$,
$h_{r,t,M}(\varphi):=2r-2r\cdot\min\left\{1,\frac{\text{dist}_{C([0,t];E)}(\varphi,K^{0}_{t}(2r))}{M}\right\},\;\varphi\in C([0,t];E).$ (4.14)
By this definition, it is clear that $h_{r,t,M}$ is bounded and Lipschitz continuous, with a Lipschitz constant of $\frac{2r}{M}$. Moreover, by the measurability of $\bar{u}^{0}_{\epsilon}$ on $C([0,t];E)$, the real valued random variable $h_{r,t,M}(\bar{u}^{0}_{\epsilon})$ is also measurable. Note that if $\text{dist}_{C([0,t];E)}(\bar{u}^{0}_{\epsilon},K^{0}_{t}(2r))\geq M$, then $h_{r,t,M}(\bar{u}^{0}_{\epsilon})=0$. Thus,
$\mathbb{P}\left(\text{dist}_{C([0,t];E)}(\bar{u}_{\epsilon}^{0},K^{0}_{t}(2r))\geq M\right)\leq\mathbb{E}\exp\left(-\frac{h_{r,t,M}(\bar{u}^{0}_{\epsilon})}{\epsilon}\right).$ (4.15)
Define by $\bar{u}^{0}_{\epsilon,f}$ the dynamics of the following equation on $E$
[…]
Then, by the variational representations in Theorem 3 of [budhiraja2008large], for any $\epsilon>0$
$\begin{split}\epsilon\log\mathbb{E}\exp\left(-\frac{h_{r,t,M}(\bar{u}^{0}_{\epsilon})}{\epsilon}\right)=-\inf_{f\in L^{2}(0,t;H)}\mathbb{E}\left\{\frac{1}{2}\int_{0}^{t}\|f(s)\|_{H}^{2}\;ds+h_{r,t,M}(\bar{u}^{0}_{\epsilon,f})\right\}.\end{split}$ (4.17)
By the definition of infimum, for any $0<\eta\leq\min\{1,r/2\}$, there exists $f_{n}\in L^{2}(0,t;H)$ such that
$\begin{split}&\ \ \ \ \mathbb{E}\left\{\frac{1}{2}\int_{0}^{t}\|f_{n}(s)\|_{H}^{2}\;ds+h_{r,t,M}(\bar{u}^{0}_{\epsilon,f_{n}})\right\}\\ &\leq\inf_{f\in L^{2}(0,t;H)}\mathbb{E}\left\{\frac{1}{2}\int_{0}^{t}\|f(s)\|_{H}^{2}\;ds+h_{r,t,M}(\bar{u}^{0}_{\epsilon,f})\right\}+\eta.\end{split}$ (4.18)
We claim that such $f_{n}$ must satisfy $\frac{1}{2}\int_{0}^{t}\|f_{n}(s)\|_{H}^{2}\;ds\leq 2r+1$. This is due to the fact that $h_{r,t,M}$ is bounded by $2r$. Indeed, if $\frac{1}{2}\int_{0}^{t}\|f_{n}(s)\|_{H}^{2}\;ds>2r+1$, then by (4.18), one gets
$\inf_{f\in L^{2}(0,t;H)}\mathbb{E}\left\{\frac{1}{2}\int_{0}^{t}\|f(s)\|_{H}^{2}\;ds+h_{r,t,M}(\bar{u}^{0}_{\epsilon,f})\right\}>2r.$
However, by taking $f=0$, one gets
$\inf_{f\in L^{2}(0,t;H)}\mathbb{E}\left\{\frac{1}{2}\int_{0}^{t}\|f(s)\|_{H}^{2}\;ds+h_{r,t,M}(\bar{u}^{0}_{\epsilon,f})\right\}\leq 2r,$
which contradicts.
Similarly, by the definition of the rate function
$\begin{split}\inf_{\varphi\in C([0,t];E)}\{I_{t}^{0}(\varphi)+h_{r,t,M}(\varphi)\}&=\inf_{f\in L^{2}(0,t;H)}\left\{\frac{1}{2}\int_{0}^{t}\|f(s)\|_{H}^{2}\;ds+h_{r,t,M}(z^{0}_{f})\right\}\\ &\leq\frac{1}{2}\int_{0}^{t}\|f_{n}(s)\|_{H}^{2}\;ds+h_{r,t,M}(z^{0}_{f_{n}}).\end{split}$ (4.19)
By substituting (4.18) and (4.19) into (4.17), we get
$\ \ \ \ \epsilon\log\mathbb{E}\exp\left(-\frac{h_{r,t,M}(\bar{u}^{0}_{\epsilon})}{\epsilon}\right)+\inf_{\varphi\in C([0,t];E)}\{I_{t}^{0}(\varphi)+h_{r,t,M}(\varphi)\}$
$\leq-\mathbb{E}\left\{\frac{1}{2}\int_{0}^{t}\|f_{n}(s)\|_{H}^{2}\;ds+h_{r,t,M}(\bar{u}^{0}_{\epsilon,f_{n}})\right\}+\frac{1}{2}\int_{0}^{t}\|f_{n}(s)\|_{H}^{2}\;ds+h_{r,t,M}(z^{0}_{f_{n}})+\eta$
$\leq\mathbb{E}\left[h_{r,t,M}(z^{0}_{f_{n}})-h_{r,t,M}(\bar{u}^{0}_{\epsilon,f_{n}})\right]+\eta$
$\leq\frac{2r}{M}\mathbb{E}\|z^{0}_{f_{n}}-\bar{u}^{0}_{\epsilon,f_{n}}\|_{C([0,t];E)}+\eta.$
Here, the last inequality is due to the Lipschitz continuity of $h_{r,t,M}$. Since $\frac{1}{2}\int_{0}^{t}\|f_{n}(s)\|_{H}^{2}\;ds\leq 2r+1$, as in the proof of Proposition 6.1 of [cerrai2003stochastic], we have that there exists $C_{r}>0$ that depends on $r$ such that for any $0<\epsilon\leq 1$,
$\sup_{t\geq 0}\|z^{0}_{f_{n}}(t)\|_{E}\leq C_{r},\;\mathbb{E}\left[\sup_{t\geq 0}\|\bar{u}^{0}_{\epsilon,f_{n}}(t)\|\right]\leq C_{r}.$
Hence, by choosing $M\geq 8C_{r}$, one gets that for any $0<\epsilon\leq 1$ and all $t\geq 0$
$\epsilon\log\mathbb{E}\exp\left(-\frac{h_{r,t,M}(\bar{u}^{0}_{\epsilon})}{\epsilon}\right)+\inf_{\varphi\in C([0,t];E)}\{I_{t}^{0}(\varphi)+h_{r,t,M}(\varphi)\}\leq\frac{4C_{r}}{M}\cdot r+\eta\leq r.$ (4.20)
Note that for any $\varphi\in C([0,t];E)$, either $\varphi\in K^{0}_{t}(2r)$ in which case $h_{r,t,M}(\varphi)=2r+1$, or $\varphi\notin K^{0}_{t}(2r)$ in which case $I_{t}^{0}(\varphi)>2r$. This implies that
$\inf_{\varphi\in C([0,t];E)}\{I_{t}^{0}(\varphi)+h_{r,t,M}(\varphi)\}\geq 2r.$ (4.21)
Now, substituting (4.20), (4.21) and (4.17) into (4.15) , we get for any $0<\epsilon\leq 1$ and all $t\geq 0$
$\ \ \ \ \epsilon\log\mathbb{P}\left(\text{dist}_{C([0,t];E)}(\bar{u}_{\epsilon}^{0},K^{0}_{t}(2r))\geq M\right)$
$\leq\epsilon\log\mathbb{E}\exp\left(-\frac{h_{r,t,M}(\bar{u}^{0}_{\epsilon})}{\epsilon}\right)+\inf_{\varphi\in C([0,t];E)}\{I_{t}^{0}(\varphi)+h_{r,t,M}(\varphi)\}-2r$
$\leq-r,$
which is the desired result.
The proof of Proposition 4.2 is complete. ∎
4.2 The LDP lower bound
@@
Fix $\zeta\in E$, $\delta>0$, $\gamma>0$, and $T>0$. We assume that $U(\zeta)<\infty$, or otherwise the result would be trivial. We set $\left\{z^{x}\right\}_{x\in E}\subset C([0,T];E)$ to be a family of paths satisfying
@@
Moreover, according to Section 2, $\left\{\mathcal{L}(\bar{u}_{\epsilon}^{x})\right\}_{\epsilon>0}$ satisfies the Freidlin-Wentzell uniform large deviations principle in $C([0,T];E)$. So, for every $r_{0},\gamma>0$ there exists $\epsilon_{0}>0$ such that $\gamma/2+\epsilon_{0}\ln\frac{1}{2}>0$ and for any $z^{x}\in C([0,T];E)$ with $I_{T}^{x}(z^{x})\leq r_{0}$,
$\inf_{x\in B_{E}(0,R)}\mathbb{P}\left(\left\|\bar{u}_{\epsilon}^{x}-z^{x}\right\|_{C([0,T];E)}<\delta/2\right)\geq\inf_{x\in B_{E}(0,R)}\exp\left(-\frac{1}{\epsilon}\left[I_{T}^{x}(z^{x})+\gamma/2+\epsilon\ln\frac{1}{2}\right]\right),$ (4.24)
for every $\epsilon\leq\epsilon_{0}$. Since the family $\left\{\bar{\mu}_{\epsilon}\right\}_{\epsilon>0}$ is the invariant measure of equation (1.3), from (4.22), (4.23) and (4.24), we have
@@
For this, denote the solution of equation (2.3) of the skeleton function by $z^{x}_{f}$. By the definition of $U(\zeta)$, there exist $T_{2}>0$, $z_{\bar{f}}^{m_{L}-\psi}\in C([0,T];E)$ and $\bar{f}\in L^{2}([0,T];H)$ such that $z^{m_{L}-\psi}_{\bar{f}}(0)=m_{L}-\psi$, $z^{m_{L}-\psi}_{\bar{f}}(T_{2})=\zeta$ and
@@
Thus, by the compactness of $\overline{B_{W^{k^{\star},p^{\star}}}(0,R)}$ in $E$ and (2.8), for any $\eta>0$ we can take $T_{1}=T_{1}(\eta)$ large enough such that
$\sup_{x\in\overline{B_{W^{k^{\star},p^{\star}}}(0,R)}}\left\|z^{x}_{0}(T_{1})-(m_{L}-\psi)\right\|_{E}\leq\eta.$
@@
$\sup_{x\in\overline{B_{W^{k^{\star},p^{\star}}}(0,R)}}\left\|z^{x}_{f}(T)-\zeta\right\|_{E}$ $\leq\sup_{y\in B_{E}(m_{L}-\psi,\lambda)}\left\|z^{y}_{\bar{f}}(T_{2})-z^{m_{L}-\psi}_{\bar{f}}(T_{2})\right\|_{E}$
$\leq\sup_{y\in B_{E}(m_{L}-\psi,\lambda)}c(T,\|y\|_{E},U(\zeta))\cdot\|y\|_{E}.$
@@
In conclusion, the path $z^{x}(t)=z^{x}_{f}(t)$ satisfies the two desired conditions. The proof is complete.
4.3 The LDP upper bound